# V2X-Real: a Largs-Scale Dataset for Vehicle-to-Everything Cooperative Perception

### v1

@@ Abstract
Recent advancements in Vehicle-to-Everything (V2X) technologies have enabled autonomous vehicles to share sensing information to see through occlusions, greatly boosting the perception capability. However, there are no real-world datasets to facilitate the real V2X cooperative perception research – existing datasets either only support Vehicle-to-Infrastructure cooperation or Vehicle-to-Vehicle cooperation. In this paper, we propose a dataset that has a mixture of multiple vehicles and smart infrastructure simultaneously to facilitate the V2X cooperative perception development with multi-modality sensing data. Our V2X-Real is collected using two connected automated vehicles and two smart infrastructures, which are all equipped with multi-modal sensors including LiDAR sensors and multi-view cameras. The whole dataset contains 33K LiDAR frames and 171K camera data with over 1.2M annotated bounding boxes of 10 categories in very challenging urban scenarios. According to the collaboration mode and ego perspective, we derive four types of datasets for Vehicle-Centric, Infrastructure-Centric, Vehicle-to-Vehicle, and Infrastructure-to-Infrastructure cooperative perception. Comprehensive multi-class multi-agent benchmarks of SOTA cooperative perception methods are provided. The V2X-Real dataset and benchmark codes will be released.
@@
1 Introduction
The evolution of autonomous driving technology has been significantly accelerated by advancements in deep learning, particularly in perception tasks crucial for safe navigation and decision-making. While progress in various perception tasks has been notable, the single-vehicle vision systems still suffer from occlusions and limited perception range. The bottlenecks mainly stem from the fact that each vehicle can only assess its surroundings from a single view point, leading to partial understanding of the environment. To address these issues, recent works [18, 26, 27, 9, 20, 13] have investigated Vehicle-to-Everything (V2X) Cooperative Perception, where each connected agent within the communication range can share the sensing information (e.g., raw LiDAR point clouds, detection outputs and intermediate neural features) with each other and leverage neighboring agent’s sensing information for visual reasoning to enlarge the perception range and see through occlusions.
Despite the great potential, there is no open V2X dataset from real-world scenarios for researchers to work with. The majority of available V2X datasets such as OPV2V [26], V2X-Sim [9] and V2X-Set [25] are generated through simulators [3] with simulated traffic dynamics and sensor renderings. This approach introduces a significant simulation-to-reality gap, thereby impeding the effective evaluation and deployment of V2X algorithms in real world [23]. Recently, DAIR-V2X [27] and V2V4Real [23] present the pioneering real-world datasets for Vehicle-to-Infrastructure (V2I) and Vehicle-to-Vehicle (V2V) collaborations. However, these datasets are limited by single collaboration mode (i.e. either V2V or V2I), involving at most two agents within the same spatial vicinity. Nevertheless, in real world, V2X collaborations can dynamically adapt to multifaceted interactions inherent in real-world traffic scenarios, including V2V, V2I, I2I and thus truly V2X. Developing algorithms based on single collaboration mode will greatly constrain the scope of V2X collaborations and hinder deploying the algorithms in complex V2X scenarios. In addition, existing dataset are all vehicle-centric by assuming vehicle as the ego agent, ignoring the pivotal role of infrastructure; however, the infrastructure perception is vital for both autonomous driving and intelligent transportation system. On one hand, it can augment the perception systems of self-driving cars by sharing sensing information with nearby agents. On the other hand, infrastructure perception can provide a reliable basis for various intelligent transportation tasks such as traffic management, traffic monitoring and signal control, which will greatly benefit from the V2I and I2I collaborations. Moreover, DAIR-V2X [27] and V2V4Real [23] are mostly collected in suburban and rural areas with an average of 11-12 objects per scene. This relatively low traffic density limits their effectiveness in testing V2X cooperative perception in more challenging scenarios and restricts the scalability of V2X collaborations.
To this end, we introduce the V2X-Real, a large-scale multi-modal multi-view dataset tailored for V2X research, including 33K LiDAR frames, 171K Camera images and over 1.2M 3D bounding box annotations. The proposed dataset is collected via two smart infrastructure and two autonomous vehicles equipped with full sensor suits (Fig. 1). To ensure the data diversity, we gather the data in two types of scenarios i.e., V2X smart intersection and V2V corridors. With over 75 hours of collected driving logs, we carefully curate and select 70 representative scenarios to form the final dataset. Unlike previous datasets, V2X-Real is exclusively collected in urban and freeway environments, featuring an average of 36 objects per scene. Such environments offer a high density of traffic and a significant presence of vulnerable road users, providing rich challenging scenarios for V2X cooperative perception research. To innovate and support a wide range of research interests in V2X cooperative perception, the dataset is divided into four specialized sub-datasets based on the type of ego agent and their collaboration modes: V2X-Real-VC for Vehicle-Centric, V2X-Real-IC for Infrastructure-Centric, V2X-Real-V2V for Vehicle-to-Vehicle, and V2X-Real-I2I for Infrastructure-to-Infrastructure cooperative perception.
@@ 1 Introduction
  • We conduct comprehensive benchmarks for multi-class multi-agent V2X cooperative perception for V2X-Real-VC, V2X-Real-IC, V2X-Real-V2V, and V2X-Real-I2I. We will release all the benchmark codes.
@@ 2.1 Self-driving datasets
Over the past few decades, a number of autonomous driving dataset has been released, greatly promoting the research and development of self-driving technologies. Among these datasets, KITTI is a pioneering autonomous driving dataset featured with full sensor suits[5]. Following KITTI, the NuScenes[1] and Waymo Open datasets[16] have expanded the scale of data significantly, offering both LiDAR point clouds and multi-view cameras that cover a 360-degree field of view, thereby enhancing the depth and breadth of research possibilities in the field. However, these datasets are all collected by single vehicle. OPV2V introduces the first V2V cooperative perception dataset, created using CARLA simulator[3] and OpenCDA[21, 24, 28]. V2XSet and V2X-Sim[9] further extended to V2X scenarios in CARLA simulator. In contrast to these simulated datasets, DAIR-V2X [27] is the first real-world dataset for V2I cooperative perception while V2V4Real [23] presents the first real-world V2V dataset collected by two connected vehicles.
Nevertheless, existing V2X datasets all focus on single collaboration mode i.e., either V2V or V2I and didn’t incorporate I2I collaborations. Furthermore, the benchmarks and evaluations within these datasets are vehicle-centric, treating infrastructure merely as an auxiliary collaborator. Infrastructure-centric collaborations such as V2I and I2I are rarely discussed, highlighting a gap in the comprehensive evaluation of V2X interactions. Compared with these datasets, our dataset features with all three V2X collaboration modes i.e., V2V, V2I, and I2I, and investigates and benchmarks both Vehicle-Centric and Infrastructure Centric V2X cooperative perception. Moreover, it is uniquely collected in congested urban environments with dense traffic and pedestrian flow, resulting in a record number of annotated 3D boxes compared to existing datasets. Furthermore, to facilitate multi-view camera-based perception, we will also provide multi-view stereo cameras covering 360 field-of-view for the vehicle agent and high-resolution network camera data for the infrastructure while only single-view cameras are provided in DAIR-V2X and V2V4Real. Additionally, existing V2X dataset only benchmark the detection performance based on single vehicle category, ignoring the perception performance for vulnerable road users. In contrast, we provide benchmarks for all the annotation types and will release our codebase for multi-class cooperative perception to facilitate V2X research on vulnerable road users, which is an important yet under-explored research area.
@@
Infra LiDAR Ouster OS1-128/64 ($\times$1)
128/64 beams, 10Hz capture frequency, $360^{\circ}$ horizontal FOV, $-16.6^{\circ}$ to $+16.6^{\circ}$(64), $-22.5^{\circ}$ to $+22.5^{\circ}$(128) vertical FOV, $\leq$ 40m range
Camera Axis-P14555 ($\times$2)
RGB, 30Hz capture frequency, 1920 $\times$ 1080 resolution, $76^{\circ}$ horizontal FOV, $48^{\circ}$ vertical FOV, JPEG compressed
GPS Garmin 18x LVC GPS ($\times$1)
5Hz update rate
Vehicle LiDAR RoboSense Ruby Plus ($\times$1)
128 beams, 10Hz capture frequency, $360^{\circ}$ horizontal FOV, $-25^{\circ}$ to $+15^{\circ}$ vertical FOV, $\leq$ 200m range
Camera ZED 2i ($\times$4)
Stereo RGBD, 10Hz capture frequency, 1920 $\times$ 1080 resolution, $120^{\circ}$ horizontal FOV, JPEG compressed
GPS/IMU Gongji GPS System ($\times$1)
1000Hz update rate, Double-Precision
@@ 2.2 V2X Cooperative Perception
Cooperative perception aims to enhance the perception capability of the V2X system by leveraging the shared information among connected agents. According to its fusion strategies, it can be broadly classified into three categories: 1) Early Fusion where the raw LiDAR point clouds are transmitted and the ego agent will aggregate the raw data to predict objects, 2) Late Fusion [14] where the detection outputs are circulated and then fused into a consistent prediction, 3) Intermediate Fusion [18, 25, 26, 6, 10, 15]where the intermediate neural features are shared and fused. Though early Fusion preserves the raw sensing information and thus can potentially reach higher accuracy, the high bandwidth requirement of transmitting raw data makes it infeasible in real-world scenario. On the other hand, Late Fusion requires minimal bandwidth, but it will lose the context information during transmission, leading to inferior performance. For a good-trade-off between bandwidth requirements and accuracy, Intermediate Fusion has increasing attentions. AttFuse [26] presents an attention-based fusion strategy to refine multi-agent’s BEV features. V2VNet [18] employs a graph neural network architecture to iteratively refine features and conduct joint perception and prediction. F-Cooper [2] presents a pooling mechanism to identify salient features. V2X-ViT [25] and CoBEVT [22] employ a sparse vision transformer for V2X collaborations. HM-ViT [20] presents a 3D heterogeneous graph transformer for camera and LiDAR fusion.
@@ 3.1 Data Acquisition
Data Collection: To ensure the diversity of our V2X data, We collected our data in two types of scenarios – V2X intersection and V2V corridor. In the intersection scenarios, two smart infrastructure units and two autonomous vehicles are operated simultaneously. Our autonomous vehicles will follow predefined routes, covering different V2X interactions. Through a combination of basic maneuvers of each autonomous vehicle at the intersection – such as left turn, right turn, waiting for traffic signal, and going straight – we designed a total of 50 distinct V2X interactions in our dataset including car following, two cars driving towards each other, two cars making left (right) turn together et al. As the V2X data is primarily collected in a single smart intersection and its surrounding streets for diverse V2X interactions, to enlarge the diversity of our dataset, we further collected separate V2V scenarios in urban corridor. In this V2V data, we designed 10 types of V2V interactions including car following, lane change, and overtaking between two autonomous vehicles et al.
Our data collection efforts span 3 months in diverse illumination conditions, with a total of over 75 hours of collected driving logs. We carefully curate the data and select 70 representative scenarios for annotation with a total of 33K LiDAR frames and 171K Camera images, providing a rich resource for V2X research.
@@
Coordinate system: There are three types of coordinate systems in our dataset, namely LiDAR coordinate, camera coordinate, and the map coordinate. Each agent maintains a separate LiDAR and camera coordinate frame. We built a dense LiDAR point cloud map [19] covering all the driving routes and leverage map matching combined with GPS/IMU positional inputs [4] to identify the transformation between each agent’s LiDAR coordinate frame and map coordinate. For 3D-2D associations, intrinsic and extrinsic calibrations between camera and LiDAR are conducted. The visualization of the calibration result is shown in Fig. 2.
Synchronization: For the multi-agent datasets, it is crucial to synchronize the time clock of all the agents. We synchronize all the computer’s time clock with GPS time to ensure consistent timestamp across sensing messages and further lock the LiDAR phase with GPS signals to ensure the temporal synchrony of multi-agents sensing observation. As the ZED2i stereo camera and Axis network camera can’t be hardware-triggered with GPS signals for synchronization, we post-process the collected data to identify the closest camera message as per LiDAR frame to reach minimal time difference. Messages with large time gap are filtered to ensure suitable synchronization between multi-agent multi-modal data.
@@ 3.2 Data Annotation and Processing
Figure 4: Polar density map of annotated bounding boxes for (a) Car, (b) Pedestrian, and (c) Truck. The polar and radial axes represent the distance and angle of the bounding boxes respectively in ego coordinate frame. The color indicates the number of box in the bin and the darker color corresponds to a larger density.
Additionally, we have implemented several strategies to enhance the efficiency and accuracy of the annotation process. We combine two infrastructure’s LiDAR point clouds to form a holistic view for annotation (Fig. 3). The combined LiDAR point clouds can efficiently reduce the annotation workload and provide dense sensing measurements for objects of interest, greatly increasing the annotation efficiency and accuracy. After annotating the 3D boxes in combined LiDAR point clouds, we will project the boxes into each infrastructure’s coordinate frame and for each box, we will count the number of LiDAR points within the box. If the number is greater than a threshold (i.e., 5 in our dataset), we will add the box to this infrastructure’s annotation result, otherwise, the box is ignored for this agent due to its sparse sensing measurements. Afterwords, we manually check the result and refine the annotations for each of the agent. For vehicles, their poses vary from scene to scene with different level of localization error, thus separate annotations for each of the vehicle are conducted to ensure accurate annotations. Additionally, for the static objects (e.g., vehicles that are parked and those awaiting the change of a traffic signal), we provide annotators with a script to automatically interpolate their poses in one frame to a sequence of frames. This greatly increased the annotation accuracy and efficiency for static objects.
@@
As the objects are annotated separately in different agents’ coordinate frames, identical objects may be assigned with different IDs. To ensure consistent object ID, we project all the annotated boxes from different agents to the same map coordinate, and calculate their IoU overlap. If their overlap is greater than a certain threshold, they are considered as the same object and thus a single unique ID is assigned.
@@
Figure 6: Bounding box LiDAR point density distribution and box size distribution for (a) Car, (b) Pedestrian and (c) Truck. In each category’s figure, the left figure depicts the number of lidar points within bounding boxes as per the radial distance from the data collection vehicle while the right figure demonstrates the bounding box size distribution.

### v2

# V2X-Real: a Large-Scale Dataset for Vehicle-to-Everything Cooperative Perception

@@ Abstract
Recent advancements in Vehicle-to-Everything (V2X) technologies have enabled autonomous vehicles to share sensing information to see through occlusions, greatly boosting the perception capability. However, there are no real-world datasets to facilitate the real V2X cooperative perception research – existing datasets either only support Vehicle-to-Infrastructure cooperation or Vehicle-to-Vehicle cooperation. In this paper, we present V2X-Real, a large-scale dataset that includes a mixture of multiple vehicles and smart infrastructure to facilitate the V2X cooperative perception development with multi-modality sensing data. Our V2X-Real is collected using two connected automated vehicles and two smart infrastructure, which are all equipped with multi-modal sensors including LiDAR sensors and multi-view cameras. The whole dataset contains 33K LiDAR frames and 171K camera data with over 1.2M annotated bounding boxes of 10 categories in very challenging urban scenarios. According to the collaboration mode and ego perspective, we derive four types of datasets for Vehicle-Centric, Infrastructure-Centric, Vehicle-to-Vehicle, and Infrastructure-to-Infrastructure cooperative perception. Comprehensive multi-class multi-agent benchmarks of SOTA cooperative perception methods are provided. The V2X-Real dataset and codebase are available at https://mobility-lab.seas.ucla.edu/v2x-real.
Keywords:
V2X Dataset Cooperative Perception Autonomous Driving
1 Introduction
The evolution of autonomous driving technology has been significantly accelerated by advancements in deep learning, particularly in perception tasks  [17, 23, 27, 9, 38] crucial for safe navigation [11, 28, 8, 39] and decision-making [24]. While progress in various perception tasks has been notable, the single-vehicle vision systems still suffer from occlusions and limited perception range. The bottlenecks mainly stem from the fact that each vehicle can only assess its surroundings from a single view point, leading to a partial understanding of the environment. To address these issues, recent works [22, 35, 36, 12, 26, 16] have investigated Vehicle-to-Everything (V2X) Cooperative Perception, where each connected agent within the communication range can share the sensing information (e.g., raw point clouds, detection results and intermediate neural features) with each other and leverage neighboring agent’s sensing information for visual reasoning to enlarge the perception range and see through occlusions.
Despite the great potential, there is no open V2X dataset from real-world scenarios for researchers to work with. The majority of available V2X datasets such as OPV2V [35], V2X-Sim [12] and V2X-Set [34] are generated through simulators [3] with simulated traffic dynamics and sensor renderings. This approach introduces a significant simulation-to-reality gap, thereby impeding the effective evaluation and deployment of V2X algorithms in the real world [32]. Recently, DAIR-V2X [36] and V2V4Real [32] present the pioneering real-world datasets for Vehicle-to-Infrastructure (V2I) and Vehicle-to-Vehicle (V2V) collaborations. However, these datasets are limited by a single collaboration mode (i.e. either V2V or V2I), involving at most two agents within the same spatial vicinity. Nevertheless, in real world, V2X collaborations can dynamically adapt to multifaceted interactions inherent in real-world traffic scenarios, including V2V, V2I, I2I and thus truly V2X. Developing algorithms based on single collaboration mode will greatly constrain the scope of V2X collaborations and hinder deploying the algorithms in complex V2X scenarios. In addition, existing datasets are vehicle-centric by assuming the vehicle as the ego agent, ignoring the pivotal role of infrastructure; however, the infrastructure perception is vital for both autonomous driving and intelligent transportation systems. On the one hand, it can augment the perception systems of self-driving cars by sharing sensing information with nearby agents. On the other hand, infrastructure perception can provide a reliable basis for various intelligent transportation tasks such as traffic management, traffic monitoring and signal control, which will greatly benefit from the V2I and I2I collaborations. Moreover, DAIR-V2X [36] and V2V4Real [32] are mostly collected in suburban and rural areas with an average of 11-12 objects per scene. This relatively low traffic density limits their effectiveness in testing V2X cooperative perception in more challenging scenarios and restricts the scalability of V2X collaborations.
@@ 1 Introduction
To this end, we introduce the V2X-Real, a large-scale multi-modal multi-view dataset tailored for V2X research, including 33K LiDAR frames, 171K camera images and over 1.2M 3D bounding box annotations. The proposed dataset is collected via two smart infrastructure and two autonomous vehicles equipped with full sensor suits (Fig. 1). To ensure the data diversity, we gather the data in two types of scenarios i.e., V2X smart intersection and V2V corridors. With over 75 hours of collected driving logs, we carefully curate and select 68 representative scenarios to form the final dataset. Unlike previous datasets, V2X-Real is primarily collected in urban environments, featuring an average of 36 objects per scene. Such environments offer a high density of traffic and a significant presence of vulnerable road users, providing rich challenging scenarios for V2X cooperative perception research. To innovate and support a wide range of research interests in V2X cooperative perception, the dataset is divided into four specialized sub-datasets based on the type of ego agent and their collaboration modes: V2X-Real-VC for Vehicle-Centric, V2X-Real-IC for Infrastructure-Centric, V2X-Real-V2V for Vehicle-to-Vehicle, and V2X-Real-I2I for Infrastructure-to-Infrastructure cooperative perception.
@@
  • We conduct comprehensive benchmarks for multi-class multi-agent V2X cooperative perception for V2X-Real-VC, V2X-Real-IC, V2X-Real-V2V, and V2X-Real-I2I.
@@ 2.1 Self-driving datasets
Over the past few decades, a number of autonomous driving datasets have been released, greatly promoting the research and development of self-driving technologies. Among these datasets, KITTI is a pioneering autonomous driving dataset featured with full sensor suits [5]. Following KITTI, the NuScenes [1] and Waymo Open datasets [20] have expanded the scale of data significantly, offering both LiDAR point clouds and multi-view cameras that cover a 360-degree field of view, thereby enhancing the depth and breadth of research possibilities in the field. However, these datasets are all collected by a single vehicle. OPV2V introduces the first V2V cooperative perception dataset, created using CARLA simulator [3] and OpenCDA [30, 33, 37]. V2XSet and V2X-Sim [12] further extended to V2X scenarios in the CARLA simulator. In contrast to these simulated datasets, DAIR-V2X [36] is the first real-world dataset for V2I cooperative perception while V2V4Real [32] presents the first real-world V2V dataset collected by two connected vehicles.
Nevertheless, existing V2X datasets all focus on a single collaboration mode. Furthermore, the benchmarks and evaluations within these datasets are vehicle-centric, treating infrastructure merely as an auxiliary collaborator. Infrastructure-centric collaborations such as V2I and I2I are rarely discussed, highlighting a gap in the comprehensive evaluation of V2X interactions. A concurrent work, RCooper [6], introduces an I2I dataset for roadside cooperative perception. Compared with these datasets, our dataset features with all three V2X collaboration modes i.e. V2V, V2I, and I2I, and investigates and benchmarks both Vehicle-Centric and Infrastructure Centric V2X cooperative perception. Moreover, it is uniquely collected in congested urban environments with dense traffic and pedestrian flow, resulting in a record number of annotated 3D boxes compared to existing datasets. Furthermore, to facilitate multi-view camera-based perception, we will also provide multi-view stereo cameras covering 360 field-of-view for the vehicle agent and high-resolution network camera data for the infrastructure. Additionally, existing V2X datasets only benchmark the detection performance based on a single vehicle category, ignoring the perception performance for vulnerable road users. In contrast, we provide benchmarks for all the annotation types and will release our codebase for multi-class cooperative perception to facilitate V2X research on vulnerable road users, which is an important yet under-explored research area.
@@
Infra LiDAR Ouster OS1-128/64 ($\times$1) 128/64 beams, 10Hz capture frequency, $360^{\circ}$ horizontal FOV, $-16.6^{\circ}$ to $+16.6^{\circ}$(64), $-22.5^{\circ}$ to $+22.5^{\circ}$(128) vertical FOV, $\leq$ 40m range
Camera Axis-P14555 ($\times$2) RGB, 30Hz capture frequency, 1920 $\times$ 1080 resolution, $76^{\circ}$ horizontal FOV, $48^{\circ}$ vertical FOV, JPEG compressed
GPS Garmin 18x LVC GPS ($\times$1) 5Hz update rate
Vehicle LiDAR RoboSense Ruby Plus ($\times$1) 128 beams, 10Hz capture frequency, $360^{\circ}$ horizontal FOV, $-25^{\circ}$ to $+15^{\circ}$ vertical FOV, $\leq$ 200m range
Camera ZED 2i ($\times$4) Stereo RGBD, 10Hz capture frequency, 1920 $\times$ 1080 resolution, $120^{\circ}$ horizontal FOV, JPEG compressed
GPS/IMU Gongji GPS System ($\times$1) 1000Hz update rate, Double Precision
@@ 2.2 V2X Cooperative Perception
Cooperative perception aims to enhance the perception capability of the V2X system by leveraging the shared information among connected agents. According to its fusion strategies, it can be broadly classified into three categories: 1) Early Fusion where the raw LiDAR point clouds are transmitted and the ego agent will aggregate the raw data to predict objects, 2) Late Fusion [18] where the detection outputs are circulated and then fused into a consistent prediction, 3) Intermediate Fusion [22, 34, 35, 7, 13, 19, 29]where the intermediate neural features are shared and fused. Though early Fusion preserves the raw information and thus can potentially reach higher accuracy, the high bandwidth requirement of transmitting raw data makes it infeasible in real-world scenarios. On the other hand, Late Fusion requires minimal bandwidth but loses the context information during transmission, resulting in inferior performance with compound errors. To strike a good balance between bandwidth requirements and accuracy, Intermediate Fusion has been gaining increasing attention. AttFuse [35] presents an attention-based fusion strategy to refine multi-agent’s BEV features. V2VNet [22] employs a graph neural network architecture for joint perception and prediction. F-Cooper [2] presents a pooling mechanism to identify salient features. V2X-ViT [34] and CoBEVT [31] employ a sparse vision transformer for V2X collaborations. HM-ViT [26] presents a 3D heterogeneous graph transformer for camera and LiDAR fusion.
@@ 3.1 Data Acquisition
Data Collection: To ensure the diversity of our V2X data, We collected our data in two types of scenarios – V2X intersection and V2V corridor. In the intersection scenarios, two smart infrastructure units and two autonomous vehicles are operated simultaneously. Our autonomous vehicles will follow predefined routes, covering different V2X interactions. Through a combination of basic maneuvers of each autonomous vehicle at the intersection – such as left turn, right turn, waiting for a traffic signal, and going straight – we designed a total of 50 distinct V2X interactions in our dataset including car following, two cars driving towards each other, two cars making left (right) turn together et al. As the V2X data is primarily collected in a single smart intersection and its surrounding streets for diverse V2X interactions, to enlarge the diversity of our dataset, we further collected separate V2V scenarios in urban corridor. In these V2V data, we designed 10 types of V2V interactions, including car following, lane change, and overtaking between two autonomous vehicles et al.
Our data collection efforts span 3 months in diverse illumination conditions, with a total of over 75 hours of collected driving logs. We carefully curate the data and select 68 representative scenarios for annotation with a total of 33K LiDAR frames and 171K Camera images, providing a rich resource for V2X research.
@@
Coordinate system: There are three types of coordinate systems in our dataset, namely the LiDAR coordinate, the camera coordinate, and the map coordinate. Each agent maintains a separate LiDAR and camera coordinate frame. We built a dense LiDAR point cloud map [25] covering all the driving routes and leverage map matching combined with GPS/IMU positional inputs [4] to identify the transformation between each agent’s LiDAR coordinate frame and map coordinate. For 3D-2D associations, intrinsic and extrinsic calibrations between the camera and LiDAR are conducted. The visualization of the calibration result is shown in Fig. 2.
Synchronization: For the multi-agent datasets, it is crucial to synchronize the time clock of all agents. We synchronize all the computer’s time clock with GPS time to ensure consistent timestamp across sensing messages and further lock the LiDAR phase with GPS signals to ensure the temporal synchrony of multi-agent sensing observation. As the ZED2i stereo camera and Axis network camera cannot be hardware-triggered with GPS signals for synchronization, we post-process the collected data to identify the closest camera message as per LiDAR frame to reach minimal time difference. Messages with a large time gap are filtered to ensure suitable synchronization between multi-agent multi-modal data.
@@ 3.2 Data Annotation and Processing
Figure 4: Polar density map of annotated bounding boxes for (a) Car, (b) Pedestrian, and (c) Truck. The polar and radial axes represent the distance and angle of the bounding boxes respectively in the ego coordinate frame. The color indicates the number of boxes in the bin and the darker color corresponds to a larger density.
Additionally, we have implemented several strategies to enhance the efficiency and accuracy of the annotation process. We combine two infrastructure’s LiDAR point clouds to form a holistic view for annotation (Fig. 3). The combined LiDAR point clouds can efficiently reduce the annotation workload and provide dense sensing measurements for objects of interest, greatly increasing the annotation efficiency and accuracy. After annotating the 3D boxes in combined LiDAR point clouds, we will project the boxes into each infrastructure’s coordinate frame and for each box, we will count the number of LiDAR points within the box. If the number is greater than a threshold (i.e., 5 in our dataset), we will add the box to this infrastructure’s annotation result, otherwise, the box is ignored for this agent due to its sparse sensing measurements. Afterwords, we manually check the result and refine the annotations for each of the agents. For vehicles, their poses vary from scene to scene with different levels of localization error, thus separate annotations for each of the vehicles are conducted to ensure accurate annotations. Additionally, for the static objects (e.g., vehicles that are parked and those awaiting the change of a traffic signal), we provide annotators with a script to automatically interpolate their poses in one frame to a sequence of frames. This greatly increased the annotation accuracy and efficiency for static objects.
@@
As the objects are annotated separately in different agents’ coordinate frames, identical objects may be assigned with different IDs. To ensure consistent object ID, we project all the annotated boxes from different agents to the same map coordinate and calculate their IoU overlap. If their overlap is greater than a certain threshold, they are considered the same object and thus a unique ID is assigned.
@@
Figure 6: Bounding box LiDAR point density distribution and box size distribution for (a) Car, (b) Pedestrian, and (c) Truck. In each category’s figure, the left figure depicts the number of lidar points within bounding boxes as per the radial distance from the data collection vehicle while the right figure demonstrates the bounding box size distribution.